Your sister just deposited $6,000 into an investment account. She believes that she will earn an annual return of 8.9 percent for the next 7 years. You believe that you will only be able to earn an annual return of 8.2 percent over the same period. How much more must you deposit today in order to have the same amount as your sister in 7 years?

A) $277.05

B) $791.77

C) $198.66

D) $258.58

E) $295.52
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Future value for my sister=6000*(1.089)^7

=10897.99(Approx)

Hence for me:

10897.99=P*(1.082)^7

P=10897.99/(1.082)^7

=6277.05(Approx)

Hence excess deposit to be made=6277.05-6000

=$277.05(Approx).